OpenStudy (anonymous):

Multiply. Show all of your work and be sure your answer is in simplest form. -7/8.-12

OpenStudy (anonymous):

(-7/8)(-12) = (-7/8)(-12/1) -12 and 8 have a common factor of 4 which can be removed. This gives: (-7/2)(-3/1) = 21/2